Can you show me to steps to solve this problem:
(9/14 divided by 3/7)x 1/2
Thanks for your help!
Amber

Answer by rwm(914)

start with
9/14 divided by 3/7
flip 3/7
9/14*7/3=3/2
3/2*1/2=3/4